OpenStudy (anonymous):

US insurers' costs for knee replacement surgery range from $17,627 to $25,462. Estimate the population variance and standard deviation in cost with 98% confidence based on a random sample of 10 persons who have had this surgery. The retail costs for the same procedure range from $40,640 to $58,702. Estimate the population variance and standard deviation in cost with 98% confidence based on a sample of 10 persons. I don't know how to get s1 and s2, please help!

OpenStudy (perl):

lets do this, we can estimate the standard deviation by using range / 4

OpenStudy (perl):

s1 = (25462- 17626) / 4 s2 = (58702-40640) / 4

OpenStudy (perl):

s1= 1959 s2 = 4515.5

OpenStudy (perl):

(df) s²/ chi²right < o² < (df) s² / chi²left

OpenStudy (perl):

degrees of freedom (df) is n-1

OpenStudy (perl):

there is a chi squared calculator here http://www.fourmilab.ch/rpkp/experiments/analysis/chiCalc.html

OpenStudy (perl):

do you see how you can use the calculator?

OpenStudy (perl):